OpenStudy (anonymous):

solve the equation for x. don not include "x=" in your answer log[4}^x=3/2

OpenStudy (anonymous):

5^x=24 x log5=log 24 x=log 24/ log 5 use your calculator x=1.97 5^1.97=23.82 ^_^ to make sure we check it hope it helps!

OpenStudy (anonymous):

don't you rewrite the logarithmic equation into an exponential equation in order to find x

OpenStudy (anonymous):

It depends... i'm not that sure but you probly can

OpenStudy (anonymous):

\[log_ba = k \iff b^k = a\] Therefore \[log_4x = \frac{3}{2} \iff4^\frac{3}{2} = x\]\[\implies x = (2^2)^\frac{3}{2}\]\[=2^3\]\[ = 8\]

myininaya (myininaya):

lol do not include x= in your answer? who says that?

OpenStudy (anonymous):

She's doing an online class where she submits the answers to a website, and that is probably the instructions for the problem..

myininaya (myininaya):

oh that makes sense lol

OpenStudy (anonymous):

Ahh!! your right nvm about my answer ...

OpenStudy (anonymous):

To tell you the truth my friend is over and she and I tired solving this... and it didnt come out right sorry... we should rethink our thoughts before we blurt out some nonsence